e^x

d/ dx[ ∫ (Int) dt] = x-In 2

2

True or false with full explanation


1
Expert's answer
2021-06-08T12:54:38-0400

"\int ^{e^x}_2lntdt=tlnt|^{e^x}_2-t|^{e^x}_2=xe^x-2ln2-e^x+2"


"\frac{d}{dx}(xe^x-2ln2-e^x+2)=e^x+xe^x-e^x=xe^x"


Answer: false